# Polycyclic codes over serial rings and their annihilator CSS construction

### v1

@@ 3 Polycyclic codes over $\mathscr{R}$
    $\eta:\mathscr{R}_{f}=\frac{\mathscr{R}[x]}{\langle f(x)\rangle}$ $\longrightarrow\frac{\mathscr{R}_{1}[x]}{\langle f_{1}(x)\rangle}\times\frac{\mathscr{R}_{2}[x]}{\langle f_{2}(x)\rangle}\times\dots\times\frac{\mathscr{R}_{k}[x]}{\langle f_{k}(x)\rangle},$ (7)
@@
    The condition for the ring $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ being a principal ideal ring in terms of the chain ring $\mathscr{R}_{i}$ and the polynomial $f_{i}(x)$ can be found in [26, Theorem 1].
@@
    If $f_{i}(x)$ is a regular basic irreducible polynomial in the ring $\mathscr{R}_{i}[x]$, then $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is a chain ring with the maximal ideal $\langle\gamma_{i}+\langle f_{i}(x)\rangle\rangle$, where $\gamma_{i}$ denotes the maximal ideal of the chain ring $\mathscr{R}_{i}$, see [11, Lemma 3.1].
Let $\mathscr{C}$ be an $f(x)$-polycyclic code over $\mathscr{R}$, we will denote the rank of the code $\mathscr{C}$ as $\operatorname{rank}_{\mathscr{R}}(\mathscr{C})$, i.e. the minimum number of generators needed to span $\mathscr{C}$ as $\mathscr{R}$-linear combinations. We will also denote the minimum Hamming distance of $\mathscr{C}$ as $d(\mathscr{C})$.
@@
  8. 8.
    $\mathscr{C}$ is a free over $\mathscr{R}$ if and only if every $\mathscr{C}_{i}$ is free over $\mathscr{R}_{i}$ for $i=\{1,\ldots,k\}$.
@@ 4 Description of polycyclic codes over $\mathscr{R}$ as $R$-submodules
($\begin{matrix}c^{0,0}_{0}&\cdots&c^{m_{1}-1,0}_{0}\\ c^{0,1}_{0}&\cdots&c^{m_{1}-1,1}_{0}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{0}&\cdots&c^{m_{1}-1,m_{2}-1}_{0}\\ \end{matrix}&\begin{matrix}c^{0,0}_{1}&\cdots&c^{m_{1}-1,0}_{1}\\ c^{0,1}_{1}&\cdots&c^{m_{1}-1,1}_{1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{1}&\cdots&c^{m_{1}-1,m_{2}-1}_{1}\\ \end{matrix}&\,\,\cdots&\begin{matrix}c^{0,0}_{n-1}&\cdots&c^{m_{1}-1,0}_{n-1}\\ c^{0,1}_{n-1}&\cdots&c^{m_{1}-1,1}_{n-1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{n-1}&\cdots&c^{m_{1}-1,m_{2}-1}_{n-1}\\ \end{matrix}\\ \end{array}\end{pmatrix}.$ (12)
@@
Let $m=m_{1}m_{2}$. A length $mn$ linear code $C$ over the chain ring $R$ with codewords given as in Equation (12) is called quasi-2-dimensional polycyclic (Q2DP) code if the columns of each subarray are closed under $t_{1}(x_{1})$-polycyclic shift and the rows of each subarray are closed under $t_{2}(x_{2})$-polycyclic shift.
@@
Similarly, the above theorem and corollary are true for any value of $s$. It is worth noting that $\{1,x_{1},x_{2},x_{1}x_{2}\}$ forms the basis for $\mathscr{R}$, and the representation matrices for these basis elements are $\{C_{t_{2}}^{0}\otimes C_{t_{1}}^{0},C_{t_{2}}^{0}\otimes C_{t_{1}},C_{t_{2}}\otimes C_{t_{1}}^{0},C_{t_{2}}\otimes C_{t_{1}}\}$, see [5, Section 6]. We notice that the matrix $M$ is constructed based on this set. Therefore, for the general case $s$, the matrix $M$ can be constructed using the representation matrices of the basis elements of $\mathscr{R}$ and $R$-module.
@@ 5 Euclidean dual of polycyclic codes over $\mathscr{R}$
where $d=f_{0}b_{0}+f_{1}b_{1}+\cdots+f_{n-1}b_{n-1}$. As usual, the Euclidean product of $x$ and $y$ in $\mathscr{R}^{n}$ is given by $\langle x,y\rangle=x\cdot y^{\text{tr}}$, where $y^{\text{tr}}$ denotes the transpose of $y$. The Euclidean dual code $\mathscr{C}^{\perp}$ of an $\mathscr{R}$-linear code $\mathscr{C}$ of length $n$ is defined as $\mathscr{C}^{\perp}:=\{x\in\mathscr{R}^{n}\mid\langle x,y\rangle=0,\,\forall y\in\mathscr{C}\}$. It is easy to see that the operator $\nu_{f(x)}$ is the adjoint operator of $\tau_{f(x)}$, that is $\langle\tau_{f(x)}(x),y\rangle=\langle x,\nu_{f(x)}(y)\rangle$. It is well-known that if a linear subspace is invariant under one operator, its Euclidean dual is invariant under the adjoint operator Therefore, we have the following theorem.
@@
[23, Theorem 3.5] The codes $\mathscr{C}$ and $\mathscr{C}^{\perp}$ are $f(x)$-polycyclic codes if and only if they are both $f(x)$-sequential codes, if and only if $\mathscr{C}$ is a constacyclic code.
@@ 6 Annihilator dual of polycyclic codes over $\mathscr{R}$
The set of vectors $\{e_{i}:1\leqslant i\leqslant n\}$, where each $e_{i}$ has a value of $1$ in the $i$-th position and 00 elsewhere, forms a basis for $\mathscr{R}^{n}$. The Gram matrix $A$ associated with this set is a matrix with entries $a_{i,j}=\langle e_{i},e_{j}\rangle_{f}$. For a linear code $\mathscr{C}$ over $\mathscr{R}$, we define the set $\mathscr{C}\cdot A:=\{cA\mid c\in\mathscr{C}\}$. Similar to Lemma 4 and Corollary 1 in [14], we have the following Lemma.
@@
    $\mathscr{C}^{\circ}=(\mathscr{C}\cdot A)^{\perp}$.
@@
    First, we claim that the annihilator product defined on each $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is non-degenerate. To prove it, note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the finite ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$, see [28, Chapter 1]. Thus, since the constant part of the polynomial $f(x)$ is invertible, the constant part of each polynomial $f_{i}(x)$ is also invertible, and therefore, by Theorem 10 the claim follows. According to the proof of the first statement in Proposition 3, the ring $\mathscr{R}[x]/\langle f(x)\rangle\cong\oplus_{i=1}^{k}\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is isomorphic to $\oplus_{i=1}^{k}\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle e_{C_{i}}$, thus $\mathscr{C}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}$. Furthermore, since the entries of the Gram matrix $A$ are in $\mathscr{R}$, by Proposition 1 yields $A=\oplus_{i=1}^{k}A_{i}e_{C_{i}}$, where $A_{i}$ is an $n\times n$ matrices over $\mathscr{R}$. Consequently, by Lemma 2, we have
    $\mathscr{C}^{\circ}=(\mathscr{C}\cdot A)^{\perp}$ $=\bigg{(}\big{(}\bigoplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}\big{)}\cdot\big{(}\bigoplus_{i=1}^{k}A_{i}e_{C_{i}}\big{)}\bigg{)}^{\perp}$
    $=\bigg{(}\bigoplus_{i=1}^{k}\big{(}\mathscr{C}_{i}\cdot A_{i}\big{)}e_{C_{i}}\bigg{)}^{\perp}$
    $=\bigoplus_{i=1}^{k}\big{(}\mathscr{C}_{i}\cdot A_{i}\big{)}^{\perp}e_{C_{i}}=\bigoplus_{i=1}^{k}\mathscr{C}_{i}^{\circ}e_{C_{i}}.$
@@
Suppose Let $f_{0}\in\mathscr{R}$ be an invertible element, $\mathscr{C}$ is an $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$ generated by a single polynomial $g(x)\in\mathscr{R}_{f}$, and the leading coefficient of $g(x)$ be a unit element in $\mathscr{R}$ and there exist $h(x)\in\mathscr{R}_{f}$ such that $f(x)=g(x)h(x)$. Then
@@
We will proof the first statement, and the other ones follow similarly. Note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$ [28, Chapter 1]. Thus, if the leading coefficient of $g(x)$ is a unit element in $\mathscr{R}$, then the leading coefficient of $g_{i}(x)$ is a unit element in $\mathscr{R}_{i}$ for all $i\in\{1,\ldots,k\}$. Therefore, by Theorem 12, each $\mathscr{C}_{i}$ is an annihilator self-dual code if and only if $f_{i}(x)=a_{i}g_{i}^{2}(x)$ for some unit element $a_{i}$ in $\mathscr{R}_{i}$. Then, the proof follows by Theorem 11. ∎
@@
Consider the ring $\mathscr{R}=R[x_{1},x_{2},x_{3}]/\langle x_{1}^{2}-1,x_{2}^{2}-1,x_{3}^{2}-1\rangle$, where $R$ is a chain ring with the residue field $\mathbb{F}_{q}$. Any element $r\in\mathscr{R}$ of the form $r=a_{1}+a_{2}x_{1}+a_{3}x_{2}+a_{4}x_{3}+a_{5}x_{1}x_{2}+a_{6}x_{2}x_{3}+a_{7}x_{1}x_{3}+a_{8}x_{1}x_{2}x_{3}$ can be uniquely expressed as $r_{1}e_{1}+r_{2}e_{2}+r_{3}e_{3}+r_{4}e_{4}+r_{5}e_{5}+r_{6}e_{6}+r_{7}e_{7}+e_{8}e_{8}$, where $r_{i}\in\mathscr{R}$ and
@@
where $t\in\mathbb{F}_{q}$ and $8t=1\mod p$. As special case, let $R=\mathbb{Z}_{9}$ and $t=2$.

### v3

@@ 3 Polycyclic codes over $\mathscr{R}$
    $\eta:\mathscr{R}_{f}=\frac{\mathscr{R}[x]}{\langle f(x)\rangle}$ $\longrightarrow\frac{\mathscr{R}_{1}[x]}{\langle f_{1}(x)\rangle}\oplus\frac{\mathscr{R}_{2}[x]}{\langle f_{2}(x)\rangle}\oplus\dots\oplus\frac{\mathscr{R}_{k}[x]}{\langle f_{k}(x)\rangle},$ (7)
@@
    The condition for the ring $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ being a principal ideal ring in terms of the chain ring $\mathscr{R}_{i}$ and the polynomial $f_{i}(x)$ can be found in [31, Theorem 1].
@@
    If $f_{i}(x)$ is a regular basic irreducible polynomial in the ring $\mathscr{R}_{i}[x]$, then $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is a chain ring with the maximal ideal $\langle\gamma_{i}+\langle f_{i}(x)\rangle\rangle$, where $\gamma_{i}$ denotes the maximal ideal of the chain ring $\mathscr{R}_{i}$, see [15, Lemma 3.1].
Let $\mathscr{C}$ be an $f(x)$-polycyclic code over $\mathscr{R}$, we will denote the rank of the code $\mathscr{C}$ as $\operatorname{rank}_{\mathscr{R}}(\mathscr{C})$, i.e. the minimum number of generators for spanning $\mathscr{C}$ as $\mathscr{R}$-linear combinations. We will also denote the minimum Hamming distance of $\mathscr{C}$ as $d(\mathscr{C})$.
@@ 4 Description of polycyclic codes over $\mathscr{R}$ as $R$-submodules
$\begin{pmatrix}\begin{array}[]{c|c|c|c}\begin{matrix}c^{0,0}_{0}&\cdots&c^{m_{1}-1,0}_{0}\\ c^{0,1}_{0}&\cdots&c^{m_{1}-1,1}_{0}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{0}&\cdots&c^{m_{1}-1,m_{2}-1}_{0}\\ \end{matrix}&\begin{matrix}c^{0,0}_{1}&\cdots&c^{m_{1}-1,0}_{1}\\ c^{0,1}_{1}&\cdots&c^{m_{1}-1,1}_{1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{1}&\cdots&c^{m_{1}-1,m_{2}-1}_{1}\\ \end{matrix}&\,\,\cdots&\begin{matrix}c^{0,0}_{n-1}&\cdots&c^{m_{1}-1,0}_{n-1}\\ c^{0,1}_{n-1}&\cdots&c^{m_{1}-1,1}_{n-1}\\ \vdots&\ddots&\vdots\\ c^{0,m_{2}-1}_{n-1}&\cdots&c^{m_{1}-1,m_{2}-1}_{n-1}\\ \end{matrix}\\ \end{array}\end{pmatrix}.$ (12)
@@
Let $m=m_{1}m_{2}$. A length $mn$ linear code $C$ over the chain ring $R$ with codewords given as in Equation (12) is a called quasi-2-dimensional polycyclic (Q2DP) code if the columns of each subarray are closed under $t_{1}(x_{1})$-polycyclic shift and the rows of each subarray are closed under $t_{2}(x_{2})$-polycyclic shift.
@@
Similarly, the above theorem and corollary are true for any value of $s$. It is worth noting that $\{1,x_{1},x_{2},x_{1}x_{2}\}$ forms the basis for $\mathscr{R}$, and the representation matrices for these basis elements are $\{C_{t_{2}}^{0}\otimes C_{t_{1}}^{0},C_{t_{2}}^{0}\otimes C_{t_{1}},C_{t_{2}}\otimes C_{t_{1}}^{0},C_{t_{2}}\otimes C_{t_{1}}\}$, see [8, Section 6]. We notice that the matrix $M$ is constructed based on this set. Therefore, for the general case $s$, the matrix $M$ can be constructed using the representation matrices of the basis elements of $\mathscr{R}$ and $R$-modules.
@@ 5 Euclidean dual of polycyclic codes over $\mathscr{R}$
where $d=f_{0}b_{0}+f_{1}b_{1}+\cdots+f_{n-1}b_{n-1}$. As usual, the Euclidean product of $x$ and $y$ in $\mathscr{R}^{n}$ is given by $\langle x,y\rangle=x\cdot y^{\text{tr}}$, where $y^{\text{tr}}$ denotes the transpose of $y$. The Euclidean dual code $\mathscr{C}^{\perp}$ of an $\mathscr{R}$-linear code $\mathscr{C}$ of length $n$ is defined as $\mathscr{C}^{\perp}:=\{x\in\mathscr{R}^{n}\mid\langle x,y\rangle=0,\,\forall y\in\mathscr{C}\}$. It is easy to see that the operator $\nu_{f(x)}$ is the adjoint operator of $\tau_{f(x)}$, that is $\langle\tau_{f(x)}(x),y\rangle=\langle x,\nu_{f(x)}(y)\rangle$. It is well-known that if a linear subspace is invariant under one operator, its Euclidean dual is invariant under the adjoint operator. Therefore, we have the following theorem.
@@
[28, Theorem 3.5] The codes $\mathscr{C}$ and $\mathscr{C}^{\perp}$ are $f(x)$-polycyclic codes if and only if they are both $f(x)$-sequential codes, if and only if $\mathscr{C}$ is a constacyclic code.
@@ 6 Annihilator dual of polycyclic codes over $\mathscr{R}$
The set of vectors $\{e_{i}:1\leqslant i\leqslant n\}$, where each $e_{i}$ has a value of $1$ in the $i$-th position and 00 elsewhere, forms a basis for $\mathscr{R}^{n}$. The Gram matrix $A$ associated with this set and the polynomial $f$ is a matrix with entries $a_{i,j}=\langle e_{i},e_{j}\rangle_{f}$. For a linear code $\mathscr{C}$ over $\mathscr{R}$, we define the set $\mathscr{C}A:=\{cA\mid c\in\mathscr{C}\}$. Similar to Lemma 4 and Corollary 1 in [18], we have the following lemma.
@@
    $\mathscr{C}^{\circ}=(\mathscr{C}A)^{\perp}$.
@@
    First, we claim that the annihilator product defined on each $\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is non-degenerate. To prove it, note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the finite ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$, see [33, Chapter 1]. Thus, since the constant part of the polynomial $f(x)$ is invertible, the constant part of each polynomial $f_{i}(x)$ is also invertible, and therefore, by Theorem 10 the claim follows. According to the proof of the first statement in Proposition 3, the ring $\mathscr{R}[x]/\langle f(x)\rangle\cong\oplus_{i=1}^{k}\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle$ is isomorphic to $\oplus_{i=1}^{k}\mathscr{R}_{i}[x]/\langle f_{i}(x)\rangle e_{C_{i}}$, thus $\mathscr{C}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}\cong\oplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}$. Furthermore, since the entries of the Gram matrix $A$ are in $\mathscr{R}$, by Proposition 1 yields $A=\oplus_{i=1}^{k}A_{i}e_{C_{i}}$, where $A_{i}$ is an $n\times n$ matrices over $\mathscr{R}$. Consequently, by Lemma 2, we have
    $\mathscr{C}^{\circ}=\left(\mathscr{C}A\right)^{\perp}$ $=\left(\bigg{(}\bigoplus_{i=1}^{k}\mathscr{C}_{i}e_{C_{i}}\bigg{)}\bigg{(}\bigoplus_{i=1}^{k}A_{i}e_{C_{i}}\bigg{)}\right)^{\perp}$
    $=\left(\bigoplus_{i=1}^{k}\bigg{(}\mathscr{C}_{i}A_{i}\bigg{)}e_{C_{i}}\right)^{\perp}$
    $=\bigoplus_{i=1}^{k}\bigg{(}\mathscr{C}_{i}A_{i}\bigg{)}^{\perp}e_{C_{i}}=\bigoplus_{i=1}^{k}\mathscr{C}_{i}^{\circ}e_{C_{i}}.$
@@
Let $f_{0}\in\mathscr{R}$ be an invertible element, $\mathscr{C}$ is an $f(x)$-polycyclic code of length $n$ over $\mathscr{R}$ generated by a single polynomial $g(x)\in\mathscr{R}_{f}$, and the leading coefficient of $g(x)$ be a unit element in $\mathscr{R}$ and there exist $h(x)\in\mathscr{R}_{f}$ such that $f(x)=g(x)h(x)$. Then
@@
We will proof the first statement, and the other ones follow similarly. Note that since $\mathscr{R}\cong\oplus_{i=1}^{k}\mathscr{R}_{i}$, the group of units of the ring $\mathscr{R}$ decomposes as $\mathscr{R}^{*}\cong\oplus_{i=1}^{k}\mathscr{R}^{*}_{i}$ [33, Chapter 1]. Thus, if the leading coefficient of $g(x)$ is a unit element in $\mathscr{R}$, then the leading coefficient of $g_{i}(x)$ is a unit element in $\mathscr{R}_{i}$ for all $i\in\{1,\ldots,k\}$. Therefore, by Theorem 12, each $\mathscr{C}_{i}$ is an annihilator self-dual code if and only if $f_{i}(x)=a_{i}g_{i}^{2}(x)$ for some unit element $a_{i}$ in $\mathscr{R}_{i}$. Then, the proof follows by Theorem 11. ∎
@@
Consider the ring $\mathscr{R}=R[x_{1},x_{2},x_{3}]/\langle x_{1}^{2}-1,x_{2}^{2}-1,x_{3}^{2}-1\rangle$, where $R$ is a chain ring with the residue field $\mathbb{F}_{q}$. With a similar approach as in [4], any element $r\in\mathscr{R}$ of the form $r=a_{1}+a_{2}x_{1}+a_{3}x_{2}+a_{4}x_{3}+a_{5}x_{1}x_{2}+a_{6}x_{2}x_{3}+a_{7}x_{1}x_{3}+a_{8}x_{1}x_{2}x_{3}$ can be uniquely expressed as $r_{1}e_{1}+r_{2}e_{2}+r_{3}e_{3}+r_{4}e_{4}+r_{5}e_{5}+r_{6}e_{6}+r_{7}e_{7}+r_{7}e_{8}$, where $r_{i}\in\mathscr{R}$ and
@@
where $t\in\mathbb{F}_{q}$ and $8t=1\mod p$. As a special case, let $R=\mathbb{Z}_{9}$ and $t=2$.